OpenStudy (anonymous):

Determine the empirical formula and the molecular formula of an organic compound containing 92.3% by mass carbon and the remainder by mass being hydrogen. Thanks!

OpenStudy (anonymous):



OpenStudy (anonymous):

do you have the Molar mass of the compound?? Without that I don't think I cannot help you find the molecular formula....

OpenStudy (anonymous):

Thanks! the molar mass wasn't provided. Is there any other way to find the molecular formula???

OpenStudy (anonymous):

i did some research and i think the molecular formula is C6H6. The relative molecular mass of C6H6 ( benzene) is 78 g/mol. :, C=(6*12) + H=(6*1) which is equal to 78 g/mol. Therefore C\[\frac{ 72 }{ 78} \times 100=92.3 % \] and H \[\frac{ 6 }{ 78 } \times 100= 7.7 %\]. The problem is How do I determine the molar mass 78 g/mol??

OpenStudy (anonymous):

to find the molecular formula you would have to multiply the empirical formula by a factor n.